OpenStudy (anonymous):

Which is equivalent to i903

OpenStudy (anonymous):

i assume this is \(i^{903}\) right? take the integer remainder when you divide 903 by 4

OpenStudy (anonymous):

you get 3 because 4 goes in to 900 evenly and the remainder is 3 so your answer is \(i^3=-i\)

OpenStudy (anonymous):

pattern is \(i^0=1,i^1=i, i^2=-1, i^3=-i, i^4=1 ...\) and ir repeats every 4